# Pushing the Accuracy Limit of Foundation Neural Network Models with Quantum Monte Carlo Forces and Path Integrals

### v1

@@ 5.4 Large Scale Quantum Dynamics Simulations with a ”beyond DFT” Foundation Neural Network Model
    In order to port Tinker-HP/Deep-HP (which is mainly written in Fortran 90) on Aurora, we resorted to OpenMP directives. This could be naturally done as we could start from a Fortran90/OpenACC version designed to run on Nvidia GPUs. This strategy allows for efficient data transfer between CPUs and Aurora GPUs as well as parallelization of the computationally intensive kernels. Additionnally, distributed parallelism through the native 3D domain decomposition of Tinker-HP[49, 1] can be directly exploited through MPI calls between Intel Xe GPUs. Contrary to our simulations on Nvidia GPUs, we resorted to an Aurora-compatible version of Pytorch[64] and not JAX for the computation of the neural network potential energy and forces. Following the same protocol described on Jean Zay (see previous section), we obtain 0.03 ns/day (mixte precision) per single adaptive sampling replica (32 nodes, 384 GPUs).As expected since this is a new portage of the code that will benefit of further optimization over the next months, the present production per replica on Aurora is lower than the numbers per replica obtained on Jean Zay which used the reference Tinker-HP implementation. However, Aurora benefits from its Exascale capabilities and from its associated extremely large availability of computing nodes.
@@
Using the STMV simulations, we open a new window on the impact of pH fluctuations on biological systems, particularly viruses, which is a critical factor influencing their structural dynamics and function. Traditional molecular dynamics (MD) approaches often simulate pH effects by manually adjusting the protonation states of ionizable residues to match the desired pH conditions or rely on specific constant-pH MD implementations[17] that are strongly depending on the force field quality. For STMV, as for any other viral particle, pH-dependent changes play a key role in the infection process, likely affecting capsid stability and RNA-capsid interactions. Simulating the pH transition from physiological to more acidic reflects conditions likely to be encountered during the early stages of viral cell entry and intracellular transport[82]. Implying the ”beyond DFT” Foundation Neural Network, we simulated the transition from pH$=$7 to pH$=$5 by explicitly introducing the calculated number of free protons into the system, based on pKa estimations. This approach allows proton exchange to occur dynamically and interactively with the molecular environment. As a proof of concept, we extracted a viral STMV capsomer (33,317 atoms) and quantitatively estimated the number of proton transfer events related to the 6-ns capsomer simulation (Fig. 4), providing a dynamic portrait of the molecular response to acidification.
@@
Extending this approach to the fully solvated STMV nucleocapsid (of $\sim$ 1 million atoms), we were able to conduct a 1.5-ns on-the-fly viral response to acidification, making a significant step forward in realistic, reactive virus modeling under varying pH conditions. Compared to the isolated capsomer, the current results reveal notable differences (Fig. 5). Specifically, we observed an increased number of protein-protein interactions within the capsid, particularly among residues that attempt to relax within the initial frames of the simulation. This behavior underscores the complexity of the full viral system. Given the large size and intricate nature of STMV, a longer simulation timescale is essential to fully capture its dynamic behavior and is currently underway on the Aurora machine.
Figure 5: Number of proton transfer events among key donor–acceptor pairs: water–protein, protein–water, and protein–protein. The protein component corresponds to the STMV capsid. The top panel corresponds to the full STMV nucleocapsid (left) and a cross-sectional view of capsid–RNA assembly with surrounding ions.
6 Conclusion
We presented an end-to-end integrated computational strategy for generating synthetic quantum chemistry datasets that significantly enhance the accuracy and predictive power of atomistic Foundation Machine Learning (ML) models. Our method pushes the boundaries of quantum chemical calculations, especially Quantum Monte Carlo (QMC) and selected Configuration Interaction (sCI), to unprecedented scales. Leveraging the power of exascale computing resources, we computed energies and forces with near-exact quantum accuracy for tens of thousands of molecular configurations, establishing a new state-of-the-art in dataset generation. In a near future, the next step for such high quality dataset is to complete it with exact Full-CI/full basis quality quantum computing results once logical qubits become available. As an Easter egg, we already included in the dataset a few ”chemically accurate” fully converged energy results obtained with our GPU-accelerated quantum emulator[92]. This will generalize in the future.
@@ 6 Conclusion
We demonstrated the practical applicability of our computational pipeline using FeNNix-Bio1, a new generation foundation neural network model trained with our high-fidelity datasets. Coupled with path integral-based Ring Polymer Molecular Dynamics (RPMD) and adaptive sampling, FeNNix-Bio1 successfully simulated condensed-phase quantum dynamics of a complex biological system comprising one million atoms with 32 RPMD beads to include nuclear quantum effects. This capability provides an unprecedented view into biochemical processes involving bond formation, proton transfers, and chemical reactions in realistic biosystems at quantum chemical accuracy. It is important to note that thanks to the generated dataset, the foundation model is not the only product that can be directly generated. Indeed, thanks to a recently proposed directed graph attention networks approach[19], the discussed FeNNol library can generate, within hours, new version of classical force fields such as GAFF (Generalized Amber Force Field)[98] that can be used to prepare initial simulation setups at limited cost. This work signifies a major step forward, bridging the gap between quantum accuracy and molecular dynamics simulations for biologically relevant scales. The ability to simulate multi-nanosecond quantum molecular dynamics on such large systems will profoundly impact our understanding of fundamental biochemical processes, potentially revolutionizing fields such as drug discovery, enzyme catalysis, and biomolecular engineering. Future work will explore even larger scales and more diverse chemical spaces, laying the foundation for general-purpose ML potentials of unprecedented accuracy and versatility towards predictive molecular modeling in Chemistry and Biology.
Acknowledgments
@@
QMC, sCI calculations, DFT computations on Aurora and some molecular dynamics simulation performed in this research used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (AB, KG, JPP). DFT Calculations on Polaris used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (RCC, JTK and YL).This research has used resources (Aurora and Polaris machines) of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. RCC, JTK and YL were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. CK and VA were supported by the Computational Chemical Sciences Program, under Award DE-SC0023382, funded by the U.S. Department of Energy, Office of Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division. This work was also made possible thanks to funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 810367), project EMC2 (JPP). We acknowledge EuroHPC Joint Undertaking for awarding the project IDs EHPC-DEV-2024D07-044 (E. P.) and EHPC-AI-2024A04-085 (E. P.) access to Leonardo at CINECA, Italy for DFT computations. The learning of the FeNNix-Bio1 model and some molecular dynamics simulations were performed at IDRIS, GENCI (Jean Zay machine, France) on grants No. A0150712052 (J.-P.P.) and grant GC010815453 (Grand Challenge H100 Jean Zay, J.-P.P.).

### v2

@@ V-D Large Scale Quantum Dynamics Simulations with a ”beyond DFT” Foundation Neural Network Model
    In perspective to run Tinker-HP/Deep-HP (mainly written in Fortran 90) on Aurora, we have updated the NVIDIA-based GPU code from OpenACC/CUDA to OpenMP/SYCL. This direct strategy preserves the initial code structure, paths the way to portability across multi-platform, minimizes re-implementation effort, and technically keeps intact the global optimization. On one hand, the translation from OpenACC/CUDA to OpenMP/SYCL implementation preserves control over data transfers between CPUs and GPUs with minor exceptions, as well as parallelization of the computationally intensive kernels ; while on the other, distributed parallelism through the native 3D domain decomposition of Tinker-HP[50, 1] can be directly exploited through MPI calls between Intel Xe GPUs. Contrary to our simulations on Nvidia GPUs, we resorted to an Aurora-compatible version of Pytorch[66] and not JAX for the computation of the neural network potential energy and forces. Following the same protocol described on Jean Zay (see previous section), we obtain 0.03 ns/day (mixed precision) per single adaptive sampling replica (32 nodes, 384 GPUs).As expected since this is a new portage of the code that will benefit of further optimization over the next months, the present production per replica on Aurora is lower than the numbers per replica obtained on Jean Zay which used the reference Tinker-HP implementation. One constraint is related to memory: Jean Zay’s H100 propose 80Gb/GPU while Aurora is adding 64Gb per Xe GPU. However, Aurora benefits from its Exascale capabilities and from its associated extremely large availability of computing nodes. Aurora can achieve a similar performance as Jean Zay with 15 replica (5760 GPUs, 480 nodes).
@@
Using the STMV simulations, we open a new window on the impact of pH fluctuations on biological systems, particularly viruses, which is a critical factor influencing their structural dynamics and function. Traditional molecular dynamics (MD) approaches often simulate pH effects by manually adjusting the protonation states of ionizable residues to match the desired pH conditions or rely on specific constant-pH MD implementations[17] that are strongly dependent on the force field quality. For STMV, as for any other viral particle, pH-dependent changes play a key role in the infection process, likely affecting capsid stability and RNA-capsid interactions. Simulating the pH transition from physiological to more acidic reflects conditions likely to be encountered during the early stages of viral cell entry and intracellular transport[84]. Implying the ”beyond DFT” Foundation Neural Network, we simulated the transition from pH$=$7 to pH$=$5 by explicitly introducing the calculated number of free protons into the system, based on pKa estimations. This approach allows proton exchange to occur dynamically and interactively with the molecular environment. As a proof of concept, we extracted a viral STMV capsomer (33,317 atoms) and quantitatively estimated the number of proton transfer events related to the 6-ns capsomer simulation (Fig. 4), providing a dynamic portrait of the molecular response to acidification.
@@
Extending this approach to the fully solvated STMV nucleocapsid (of $\sim$ 1 million atoms), we were able to conduct a 1.5-ns on-the-fly viral response to acidification, making a significant step forward in realistic, reactive virus modeling under varying pH conditions. Compared to the isolated capsomer, the current results reveal notable differences (Fig. 5). Specifically, we observed an increased number of protein-protein interactions within the capsid, particularly among residues that attempt to relax within the initial frames of the simulation. This behavior underscores the complexity of the full viral system. Additionally, the numerous water molecules present in the system promote frequent proton transfer events within the bulk. Given the large size and intricate nature of STMV, a longer simulation timescale is essential to fully capture its dynamical behavior and is currently underway on the Aurora machine.
Figure 5: Number of proton transfer events among key donor–acceptor pairs: water–protein, protein–water, and protein–protein. The protein component corresponds to the STMV capsid. The top panel corresponds to the full STMV nucleocapsid (left) and a cross-sectional view of capsid–RNA assembly with surrounding ions (right). The nucleocapsid images are generated with the VTX molecular visualization software[59].
VI Conclusion
We presented an end-to-end integrated computational strategy for generating synthetic quantum chemistry datasets that significantly enhance the accuracy and predictive power of atomistic Foundation Machine Learning models. Our method pushes the boundaries of quantum chemical calculations, especially Diffusion Quantum Monte Carlo and selected Configuration Interaction, to unprecedented scales. Leveraging the power of exascale computing resources, we computed energies and forces with near-exact quantum accuracy at the complete basis set limit for tens of thousands of molecular configurations, establishing a new state-of-the-art in dataset generation. In a near future, going beyond DMC and the combination of multi-determinant QMC energies and forces with selected-CI wavefunction could be difficult with classical computers. However, one next step for constructing such high accuracy dataset could be to complement it with exact Full-CI/full basis quality quantum computing results (see Figure 1) once logical qubits become available. As an Easter egg, we already included in the dataset a dozen ”chemically accurate” converged energy datapoints (i.e. equivalent to a (sCI+PT2),full basis result) obtained with our GPU-accelerated quantum emulator and extracted from a recent study[95]. This should generalize in the future.
@@ VI Conclusion
We demonstrated the practical applicability of our computational pipeline using FeNNix-Bio1, a new generation foundation neural network model trained with our high-fidelity datasets including DMC results. Coupled with path integral-based Ring Polymer Molecular Dynamics (RPMD) and adaptive sampling, FeNNix-Bio1 successfully simulated condensed-phase quantum dynamics of a complex biological system comprising one million atoms with 32 RPMD beads to fully include nuclear quantum effects. This capability provides an unprecedented view into biochemical processes involving bond formation, proton transfers, and chemical reactions in realistic biosystems at quantum chemical accuracy. It is important to note that thanks to the generated dataset, the foundation model is not the only product that can be directly generated. Indeed, thanks to a recently proposed directed graph attention networks approach[19], the discussed FeNNol library can generate, within hours, new version of classical force fields such as GAFF (Generalized Amber Force Field)[101] that can be used to prepare initial simulation setups at very limited computational cost. This work signifies a major step forward, bridging the gap between chemically accurate quantum methods and molecular dynamics simulations for biologically relevant scales. The ability to simulate multi-nanosecond quantum molecular dynamics on such large systems including their chemical reactivity thanks to a Foundation ML Model will profoundly impact our understanding of fundamental biochemical processes, potentially revolutionizing fields such as drug discovery, enzyme catalysis, and biomolecular engineering. Future work will explore even larger scales and more diverse chemical spaces, laying the foundation for general-purpose ML potentials of unprecedented accuracy and versatility towards predictive molecular modeling in Chemistry and Biology.
Acknowledgment
@@
QMC, sCI calculations, DFT computations on Aurora and some molecular dynamics simulation performed in this research used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (AB, KG, JPP). DFT Calculations on Polaris used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (RCC, JTK and YL).This research has used resources (Aurora and Polaris machines) of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. RCC (QMC forces algorithms), JTK (QMC workflow engine) and YL (QMCPACK code development) were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. CK and VA were supported by the Computational Chemical Sciences Program, under Award DE-SC0023382, funded by the U.S. Department of Energy, Office of Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division. This work was also made possible thanks to funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 810367), project EMC2 (JPP). We acknowledge EuroHPC Joint Undertaking for awarding the project IDs EHPC-DEV-2024D07-044 (E. P.) and EHPC-AI-2024A04-085 (E. P.) access to Leonardo at CINECA, Italy for DFT computations. The learning of the FeNNix-Bio1 model and some molecular dynamics simulations were performed at IDRIS, GENCI (Jean Zay machine, France) on grants No. A0150712052 (J.-P.P.) and grant GC010815453 (Grand Challenge H100 Jean Zay, J.-P.P.).
This manuscript has been authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up,irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).